# Simultaneous spatial-parametric collocation approximation for parametric PDEs with log-normal random inputs

### v3

@@ 2.2 Extended least squares sampling algorithms in Bochner spaces
$\sup_{k\in{\mathbb{N}}_{0}}2^{\alpha k}\left\|{\delta_{k}v}\right\|_{H_{X,{\boldsymbol{\sigma}}_{2}}^{\alpha,\tau}}\ \leq\ 2\|v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}^{\alpha,\tau}}\ <\infty.$
@@
Assume that there is a constant $K$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}$. Given arbitrary sample points ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{k}\in U$ and functions $h_{1},\ldots,h_{k}\in L_{2}(U,{\mathbb{C}};\mu)$, for the sampling algorithm $S_{n}^{X}$ in $L_{2}(U,X;\mu)$ defined by (2.7), we have for every $m\in{\mathbb{N}}_{0}$,
@@
$S_{n}^{{\mathbb{C}}}f:=S_{n}^{{\mathbb{C}}}({\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{n},\omega_{1},\dots,\omega_{n},\Phi_{m})f$
of a function $f\colon U\to{\mathbb{C}}$ is given by
@@
The following result [25, Theorem 3] (see also its proof) gives a error bound of the approximation by $S_{n}^{\mathbb{C}}$ via Kolmogorov $n$-widths.
@@
$S_{\Lambda}v:=\sum_{s\in\Lambda}(\delta_{m}v)_{s}\varphi_{s}.$
@@
$\Lambda_{m}(\xi):=\ \big\{\sigma_{2;s}\leq\xi^{1/q_{1}}2^{-\alpha m}m^{\tau},\ \sigma_{1;s}\leq\xi^{1/q_{1}}\big\}.$
@@
$(\delta_{m}v)_{\xi}:=\sum_{\sigma_{1;s}^{q_{1}}\leq\xi}(\delta_{m}v)_{s}\varphi_{s},$
@@
$\begin{split}\|(\delta_{m}v)-(\delta_{m}v)_{\xi}\|_{L_{2}(U,X;\mu)}^{2}\ &=\ \sum_{\sigma_{1;s}>\xi^{1/q_{1}}}\|(\delta_{m}v)_{s}\|_{X}^{2}\ =\ \sum_{\sigma_{1;s}>\xi^{1/q_{1}}}(\sigma_{1;s}\|(\delta_{m}v)_{s}\|_{X})^{2}\sigma_{1;s}^{-2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{\sigma_{1;s}>\xi^{1/q_{1}}}(\sigma_{1;s}\|(\delta_{m}v)_{s}\|_{X})^{2}\\[6.45831pt] \ &\leq\ (2K)^{2}\xi^{-2/q_{1}}\|\delta_{m}v\|_{H_{X}^{\alpha,\tau}}^{2}.\end{split}$
@@
$\begin{split}\|(\delta_{m}v)_{\xi}-S_{\Lambda_{m}(\xi)}(\delta_{m}v)\|_{L_{2}(U,X;\mu)}^{2}\ &=\ \sum_{\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}\left\|{(\delta_{m}v)_{s}}\right\|_{X}^{2}\\ &\ \leq\ \sum_{\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}\left(2^{\alpha m}m^{-\tau}\xi^{-1/q_{1}}\sigma_{2;s}\left\|{(\delta_{m}v)_{s}}\right\|_{X}\right)^{2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{s\in{\mathbb{N}}}\left(2^{\alpha m}m^{-\tau}\sigma_{2;s}\left\|{(\delta_{m}v)_{s}}\right\|_{X}\right)^{2}\\[6.45831pt] \ &\leq\ (2K)^{2}\xi^{-2/q_{1}}\|\delta_{m}v\|_{H_{X}^{\alpha,\tau}}^{2}.\end{split}$
@@
$\begin{split}{\color[rgb]{0,0,0}{\ |\Lambda_{m}(\xi)|}}\ &=\ \sum_{s\in\Lambda_{m}(\xi)}1\ =\ 2^{-m}\sum_{\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}2^{m}\\ \ &\leq\ 2^{-m}\sum_{s\in{\mathbb{N}}:\ \sigma_{1;s}\leq\xi^{1/q_{1}}}\ \sum_{(k,s)\in{\mathbb{N}}_{0}\times{\mathbb{N}}:\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha k}k^{\tau}}2^{k}\\ \ &\leq\ 2^{-m}\sum_{\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\sigma_{2;{\boldsymbol{s}}}^{-1/\alpha}\\ \ &\leq\ 2^{-m}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/q}\left(\sum_{\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}1\right)^{1/q^{\prime}}\\ \ &\leq 2^{-m}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/q}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\xi\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\right)^{1/q^{\prime}}\\ \ &\leq 2^{-m}\xi^{1+\delta/\alpha}\left(\log\xi\right)^{\tau/\alpha}.\end{split}$
@@
$\begin{split}\xi_{n}^{-1/q_{1}}\ &\leq\ C\,\left(2^{m}n(\log n+m)^{-\tau/\alpha}\right)^{-\frac{1}{q_{1}}\frac{\alpha}{\alpha+\delta}}\\ &\ \leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha}.\end{split}$
@@ 2.3 Multi-level least squares sampling algorithms
Since $\delta_{k}v$ belongs to the subspace $V_{k}+V_{k-1}$ of dimension at most $2^{k}+2^{k-1}$ and the number of sampling points in the sampling algorithms $S_{n_{k}}^{X}$ is $n_{k}$. The computational complexity $\operatorname{Comp}\left({\mathcal{S}}_{n}^{X}\right)$ of the operator ${\mathcal{S}}_{n}^{X}$ can be defined as
@@
Assume that there is a constant $K:=K_{\tau}$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in U$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{n}$, such that for the operator ${\mathcal{S}}_{n}^{X}$ defined as in (2.26)–(2.28),
@@
By Lemma 2.7, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in U$, $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$ such that
@@
Let $0<q_{1}\leq q_{2}<2$ and $\varepsilon$ be any positive number. Let the number $\beta$ be defined as in (2.22). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.29). Assume that there is a constant $K_{\varepsilon}$ such that for every $v\in H_{X}^{\alpha,1+\varepsilon}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,1+\varepsilon}}\leq K_{\varepsilon}\left\|{f}\right\|_{H_{X}^{\alpha,1+\varepsilon}}$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in U$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.26)–(2.28),
@@
Notice that if $\|{\boldsymbol{\sigma}}^{-1}\|_{\ell_{q}({\mathbb{F}})}<\infty$ for some $0<q\leq 2$, then for every $v\in B_{X,{\boldsymbol{\sigma}}}({\mathbb{U}}^{\infty})$, the series (2.34) converges absolutely and unconditionally in $L_{2}({\mathbb{U}}^{\infty},X;\gamma)$ to $v$ (see [23, Lemma 3.1] for the case $\mathbb{U}$$^{\infty}$$\mathbb{R}$$^{\infty}$, the case ${\mathbb{U}}^{\infty}={\mathbb{I}}^{\infty}$ can be proven by the same arguments). Hence, we can reorder the countable set $\mathbb{F}$} as ${\mathbb{F}}=({\boldsymbol{s}}_{j})_{j\in{\mathbb{N}}}$ so that the sequence ${\boldsymbol{\sigma}}=(\sigma_{{\boldsymbol{s}}_{j}})_{j\in{\mathbb{N}}}$ is non-decreasing. Put $U:={\mathbb{U}}^{\infty}$, $\sigma_{j}:=\sigma_{{\boldsymbol{s}}_{j}}$, $\varphi_{j}:=\phi_{{\boldsymbol{s}}_{j}}$ and $v_{j}:=v_{{\boldsymbol{s}}_{j}}$. Then $B_{X,{\boldsymbol{\sigma}}}({\mathbb{U}}^{\infty})$ can be seen as the set $B_{X,{\boldsymbol{\sigma}}}$ of all functions $v\in L_{2}(U,X;\gamma))$ represented by the series
@@
Let the assumptions of Theorem 2.2 hold for the space $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\mathbb{R}}^{\infty}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.26)–(2.28),

### v4

@@ 2.2 Extended least squares sampling algorithms in Bochner spaces
${\color[rgb]{0,0,0}{\left(\sum_{k\in{\mathbb{N}}_{0}}\left(2^{\alpha k}k^{-\tau}\left\|{\delta_{k}v}\right\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}\right)^{2}\right)^{1/2}}}\ \leq\ 2\|v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}^{\alpha,\tau}}\ <\infty.$ (2.13)
@@
[…]
@@
$\color[rgb]{0,0,0}{S_{n}^{{\mathbb{C}}}f:=S_{n}^{{\mathbb{C}}}({\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{n},\omega_{1},\dots,\omega_{n},\Phi_{m})f}$}
of a function $f\colon U\to{\mathbb{C}}$, is given by
@@
The following result [24, Theorem 3] (see also its proof) gives a error bound of the approximation by $S_{n}^{\mathbb{C}}$ via Kolmogorov $n$-widths.
@@
$S_{\Lambda}v:=\sum_{s\in\Lambda}{\color[rgb]{0,0,0}{v_{s}}}\varphi_{s}.$
@@
$\Lambda_{m}(\xi):=\ \big\{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\ \sigma_{2;s}\leq\xi^{1/q_{1}}2^{-\alpha m}m^{\tau},\ \sigma_{1;s}\leq\xi^{1/q_{1}}\big\}.$
@@
$(\delta_{m}v)_{\xi}:=\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}^{q_{1}}\leq\xi}(\delta_{m}v)_{s}\varphi_{s},$
@@
$\begin{split}\|(\delta_{m}v)-(\delta_{m}v)_{\xi}\|_{L_{2}(U,X;\mu)}^{2}\ &=\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}>\xi^{1/q_{1}}}\|(\delta_{m}v)_{s}\|_{X}^{2}\ =\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}>\xi^{1/q_{1}}}(\sigma_{1;s}\|(\delta_{m}v)_{s}\|_{X})^{2}\sigma_{1;s}^{-2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}>\xi^{1/q_{1}}}(\sigma_{1;s}\|(\delta_{m}v)_{s}\|_{X})^{2}\\[6.45831pt] \ &\leq\ (2K)^{2}\xi^{-2/q_{1}}\|\delta_{m}v\|_{H_{X}^{\alpha,\tau}}^{2}.\end{split}$
@@
$\begin{split}\|(\delta_{m}v)_{\xi}-S_{\Lambda_{m}(\xi)}(\delta_{m}v)\|_{L_{2}(U,X;\mu)}^{2}\ &=\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}\left\|{(\delta_{m}v)_{s}}\right\|_{X}^{2}\\ &\ \leq\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}\left(2^{\alpha m}m^{-\tau}\xi^{-1/q_{1}}\sigma_{2;s}\left\|{(\delta_{m}v)_{s}}\right\|_{X}\right)^{2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{s\in{\mathbb{N}}}\left(2^{\alpha m}m^{-\tau}\sigma_{2;s}\left\|{(\delta_{m}v)_{s}}\right\|_{X}\right)^{2}\\[6.45831pt] \ &\leq\ (2K)^{2}\xi^{-2/q_{1}}\|\delta_{m}v\|_{H_{X}^{\alpha,\tau}}^{2}.\end{split}$
@@
$\begin{split}{\color[rgb]{0,0,0}{\ |\Lambda_{m}(\xi)|}}\ &=\ \sum_{s\in\Lambda_{m}(\xi)}1\ =\ 2^{-m}\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}2^{m}\\ \ &\leq\ 2^{-m}\sum_{s\in{\mathbb{N}}:\ \sigma_{1;s}\leq\xi^{1/q_{1}}}\ \sum_{{\color[rgb]{0,0,0}{k\in{\mathbb{N}}_{0}:}}\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha k}k^{\tau}}2^{k}\\ \ &\leq\ 2^{-m}\sum_{\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\sigma_{2;{\boldsymbol{s}}}^{-1/\alpha}\\ \ &\leq\ 2^{-m}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\left(\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/q}\left(\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}1\right)^{1/q^{\prime}}\\ \ &\leq 2^{-m}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/q}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\xi\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\right)^{1/q^{\prime}}\\ \ &\leq 2^{-m}\xi^{1+\delta/\alpha}\left(\log\xi\right)^{\tau/\alpha}.\end{split}$
@@
$\begin{split}\xi_{n}^{-1/q_{1}}\ \leq\ C\,\left(2^{m}n(\log n+m)^{-\tau/\alpha}\right)^{-\frac{1}{q_{1}}\frac{\alpha}{\alpha+\delta}}\ \leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha}.\end{split}$
@@ 2.3 Multi-level least squares sampling algorithms
Since $\delta_{k}v({\boldsymbol{y}}_{k,i})$ belongs to the subspace $V_{k}+V_{k-1}$ of dimension at most $2^{k}+2^{k-1}$, and the number of sample points in the algorithms $S_{n_{k}}^{X}$ is $n_{k}$, the computational complexity $\operatorname{Comp}\left({\mathcal{S}}_{n}^{X}\right)$ of the operator ${\mathcal{S}}_{n}^{X}$ can be defined as
@@
Assume that there is a constant $K:=K_{\tau}$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{U_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{n}$, such that for the operator ${\mathcal{S}}_{n}^{X}$ defined as in (2.29)–(2.31),
@@
By Lemma 2.8, there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{U_{0}}}$, $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$ such that
@@
Let $0<q_{1}\leq q_{2}<2$ and $\varepsilon$ be any positive number. Let the number $\beta$ be defined as in (2.25). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.32). Assume that there is a constant $K_{\varepsilon}$ such that for every $v\in H_{X}^{\alpha,1+\varepsilon}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,1+\varepsilon}}\leq K_{\varepsilon}\left\|{f}\right\|_{H_{X}^{\alpha,1+\varepsilon}}$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{U_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.29)–(2.31),
@@
Notice that if $\|{\boldsymbol{\sigma}}^{-1}\|_{\ell_{q}({\mathbb{F}})}<\infty$ for some $0<q\leq 2$, then for every $v\in B_{X,{\boldsymbol{\sigma}}}({\mathbb{U}}^{\infty})$, the series (2.37) converges absolutely and unconditionally in $L_{2}({\mathbb{U}}^{\infty},X;\gamma)$ to $v$ (see [22, Lemma 3.1] for the case $\mathbb{U}$$^{\infty}$$\mathbb{R}$$^{\infty}$, the case ${\mathbb{U}}^{\infty}={\mathbb{I}}^{\infty}$ can be proven by the same arguments). Hence, we can reorder the countable set $\mathbb{F}$} as ${\mathbb{F}}=({\boldsymbol{s}}_{j})_{j\in{\mathbb{N}}}$ so that the sequence ${\boldsymbol{\sigma}}=(\sigma_{{\boldsymbol{s}}_{j}})_{j\in{\mathbb{N}}}$ is non-decreasing. Put $U:={\mathbb{U}}^{\infty}$, $\sigma_{j}:=\sigma_{{\boldsymbol{s}}_{j}}$, $\varphi_{j}:=\phi_{{\boldsymbol{s}}_{j}}$ and $v_{j}:=v_{{\boldsymbol{s}}_{j}}$. Then $B_{X,{\boldsymbol{\sigma}}}({\mathbb{U}}^{\infty})$ can be seen as the set $B_{X,{\boldsymbol{\sigma}}}$ of all functions $v\in L_{2}(U,X;\gamma))$ represented by the series
@@
Let the assumptions of Theorem 2.2 hold for the space $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.29)–(2.31),